six moles of nh3 was enclosed in a container of 2 litre at 25 degree celcius at equilibrium 40 percent compound was decomposed. find kc
1
Expert's answer
2019-01-28T05:39:55-0500

C(NH3)= 6mol/2L = 3 mol/L

3*0.4=1.2 mol/L NH3 was decomposed

2NH3 = N2 + 3H2

At equilibrium:

[NH3]=3-1.2=1.8 mol/L

[N2]=0.6 mol/L

[H2]=1.8 mol/L

Kc = [N2]•([H2]^3)/[NH3]^2 = 1.08